OpenStudy (anonymous):

Let f(x) = 3x + 2 and g(x) =7x + 6. Find f ·g and its domain. A. 6x2 + 4x + 42; all real numbers B. 6x2 + 4x + 42; all real numbers except x = -2/3 C. 21x2 + 32x + 12; all real numbers D. 21x2 +32x + 12; all real numbers except x = -6/7

myininaya (myininaya):

do you know the domain of f?

OpenStudy (anonymous):

idk what a domain is

myininaya (myininaya):

The domain is where the relation (or function in this case) exists for the x values.

myininaya (myininaya):

Like is there any x value you can think of that you can replace the x in 3x+2 with that would make 3x+2 not exist?

OpenStudy (anonymous):

i dont understand

myininaya (myininaya):

f(x)=3x+2 f(-3)=3(-3)+2=-9+2=-7 . -7 exists on the real number line f(0)=3(0)+2=0+2=2 . 2 exists on the real number line f(5)=3(5)+2=15+2=17 . 17 exists on the real number line can you think of a number you can plug in for x that would not give you a real number?

myininaya (myininaya):

Do you know how to graph f(x)=3x+2?

myininaya (myininaya):

Ok how about if I told you we have a function given by the following set {(3,4),(5,2),(8,9),(12,8)} Could you tell me the domain of that set?